# On a sequence of Kimberling and its relationship to the Tribonacci word

### v3

@@ Abstract
In 2017, Clark Kimberling defined an interesting sequence ${\bf B}=0100101100\cdots$ of 0’s and $1$’s by certain inflation rules, and he made a number of conjectures about this sequence and some related ones. In this note we prove his conjectures using, in part, the Walnut theorem-prover. We show how his word is related to the infinite Tribonacci word, and we determine both the subword complexity and critical exponent of $\bf B$.
@@ 1 Introduction
In June 2017, Clark Kimberling defined sequence A288462 in the OEIS [17] as follows: it is the infinite fixed point of the inflation rules $00\rightarrow 0101$, $1\rightarrow 10$, starting with $00$. In this note we prove his conjectures using, in part, the Walnut theorem-prover. We also show how his word is related to the infinite Tribonacci word, and we determine both the subword complexity and critical exponent of $\bf B$.
@@
As stated this description is perhaps slightly vague, so here is some elaboration. We start with $B_{0}=00$. To find $B_{i+1}$ from $B_{i}$, we do the following: we factor $B_{i}$ into maximal blocks of the form $00$, 0, and $1$; here maximal means we cannot extend a block further to the right or left. Then $B_{i+1}$ is the result of applying the inflation rules $0\rightarrow 0$, $1\rightarrow 10$, and $00\rightarrow 0101$ to $B_{i}$.
@@
Define $\beta_{i}=|B_{i}|$, the length of the $i$’th iterate. Let $c_{0}=2$, $c_{1}=4$, $c_{2}=6$, $c_{3}=10$, and define $c_{i}=2c_{i-1}-c_{i-4}$ for $i\geq 4$; this is sequence A288465 in the OEIS. Kimberling conjectured that $\beta_{i}=c_{i}$ for all $i\geq 0$. In this note, we prove Kimberling’s conjecture, as well as his conjectures about the related sequences A288463 and A288464. We also find other properties of the sequence $\bf B$} that link it to the infinite Tribonacci word A080843.
Kimberling indexed the sequence $\bf B$} starting at position $1$. However, for using Walnut, it is easier to index starting at position 0, and this is the convention we use in this paper at the beginning. Later on we will have to use his indexing.
@@ 2 Kimberling’s conjecture
Clearly each $1$ in a $B_{i}$ gives rise to $10$ in $B_{i+1}$, which adds $1$ to the length for each $1$ that appears. Similarly, each $00$ in a $B_{i}$ gives rise to $0101$, which adds $2$ to the length for each $00$ that appears. Defining $N_{w}(i)$ to be the number of occurrences of a word $w$ in $B_{i}$, we therefore have
$\beta_{i+1}=\beta_{i}+N_{1}(i)+2N_{00}(i).$ (1)
Two iterations of the inflation rules turn $11$ into $1001001$, $101$ into $100101100$, and $1001$ into $10010110010100$. Hence
$N_{00}(i+1)=N_{1}(i-1)+N_{00}(i-1).$ (2)
On the other hand the rule $00\rightarrow 0101$ gives
$N_{1}(i+1)=N_{1}(i)+2N_{00}(i).$ (3)
Let $T_{n}$ be the $n$’th Tribonacci number, defined by $T_{0}=0$, $T_{1}=1$, $T_{2}=1$, and $T_{n}=T_{n-1}+T_{n-2}+T_{n-3}$ for $n\geq 3$.
@@
We have
$\beta_{i}$ $=T_{i+2}+T_{i}+1$
$N_{00}(i)$ $=T_{i-1}+T_{i-2}$
$N_{1}(i)$ $=2T_{i}$
@@
Proof.
By induction on ii. The base cases are easily checked. Now assume the result is true for $i$ and we prove it for $i+1$.
From Eq. (1) we have
$\beta_{i+1}$ $=\beta_{i}+N_{1}(i)+2N_{00}(i)$
$=T_{i+2}+T_{i}+1+2T_{i}+2(T_{i-1}+T_{i-2})$
$=T_{i+3}+T_{i+1}+1.$
From Eq. (2) we have
$N_{00}(i+1)=N_{1}(i-1)+N_{00}(i-1)=2T_{i-1}+T_{i-2}+T_{i-3}=T_{i}+T_{i-1}.$
From Eq. (3) we have
$N_{1}(i+1)=N_{1}(i)+2N_{00}(i)=2T_{i}+2T_{i-1}+2T_{i-2}=2T_{i+1}.$
The proof is complete. ∎
@@ 3 Relationship to the Tribonacci word
The infinite sequence $\bf B$ given in A288462 is quite closely related to the celebrated Tribonacci word ${\bf TR}=t_{0}t_{1}t_{2}\cdots=01020100102010102010010201020100102010102010\cdots$, the fixed point of the morphism $0\rightarrow 01$, $1\rightarrow 02$, $2\rightarrow 0$. For more information about $\bf TR$, see, for example, [3].
@@
Kimberling’s sequence $\bf B$} is equal to $0f(\bf TR)$, where $f:0\to 10,\ 1\to 0,\ 2\to 1$.
@@
Let $c_{i}(n)$ denote the number of occurrences of the letter $i$ in the length-$n$ prefix of $\bf TR$. In [15, §10.12], it is shown how to obtain synchronized automata computing the maps $n\rightarrow c_{i}(n)$. By “synchronized” we mean that these automata (called c0, c1, c2) take two inputs in parallel, $n$ and $x$, in Tribonacci representation and accept if and only if $x=c_{i}(n)$. See [14] for more aboout the notion of synchronized automata.
@@
We can now write Walnut code that implements this algorithm. For more about Walnut and its use in combinatorics on words, see [15]. The first 7 lines are taken from [15, §10.12].
@@ 4 Related sequences
Similarly, we can compute $I_{0}(n)$, the index of the $n$’th 0 in $\bf B$, again with $\bf B$ indexed starting at $1$.
@@
for $n\geq 1$ and some constant $\psi\doteq 1.83$. It turns out that $\psi=1.8392867552\cdots$ is the unique real zero of the polynomial $X^{3}-X^{2}-X-1$.
@@
    $\lfloor\gamma n\rfloor-1\leq I_{1}(n)\leq\lfloor\gamma n\rfloor+2$.
@@
    Now we show that $1\leq\lfloor\psi n\rfloor-I_{0}(n)\leq 1$:
@@
    For $A_{1}(n)$ we use the code tribb from [15, §10.12]:
@@
    which is equivalent to
    $3/2\leq I_{1}(n)-\lfloor\gamma n\rfloor\leq 5/2$
@@
5 Subword complexity and critical exponent
We use the term “subword” and “factor” to mean the same concept: a contiguous block of letters within a word.
Recall that the subword complexity (aka factor complexity) of a sequence is the function mapping $n$ to the number of distinct blocks of length $n$ appearing in it. A word is called a Rote word [13] if its subword complexity is $2n$ for $n\geq 1$.
@@ 5.1 Bispecial factors of the sequence $\bf B$
In order to determine both the subword complexity and the critical exponent of $\bf B$, the knowledge of bispecial factors in $\bf B$ is essential. The description of bispecial factors and their return words in $\bf TR$ is taken from [6, 8, 7]. The sequence $(b_{n})_{n=0}^{\infty}$ of all non-empty bispecial factors, ordered by length, in the Tribonacci word, satisfies $b_{0}=0$ and for $n\geq 1$,
$b_{n}=\varphi(b_{n-1})0,\quad\text{where}\ \varphi:0\to 01,\ 1\to 02,\ 2\to 0\,.$
@@
By Theorem 2, Kimberling’sequence is equal to $f({\bf TR})$. As $\bf TR$} is fixed by $\varphi$, $\bf B$} is also the image of the Tribonacci sequence under the morphism $\pi=f\circ\varphi$. Instead of $f$ we will work with the morphism $\pi:0\to 100,\ 1\to 101,\ 2\to 10$ because it allows us to derive the form of bispecial factors in $\bf B$ in an easier way.
@@
5.2 Subword complexity of the sequence $\bf B$
@@
By Lemma 9, for each length $n\geq 1$, there are two distinct left special factors in $\bf B$, which confirms that the subword complexity is $2n$ for all $n\geq 1$. ∎
@@ 5.3 Return words to bispecial factors in $\bf B$
    The claim (b) immediately follows from (a).
@@
    $|v|\geq\min\{|\pi(\varphi^{n}(00))|,|\pi(\varphi^{n}(01))|,|\pi(\varphi^{n}(02))|\}=|\pi(\varphi^{n}(02))|=|\pi(\varphi^{n}(0))|+|\pi(\varphi^{n-1}(0))|.$

### v4

@@ Abstract
In 2017, Clark Kimberling defined an interesting sequence ${\bf B}=0100101100\cdots$ of 0’s and $1$’s by certain inflation rules, and he made a number of conjectures about this sequence and some related ones. In this note we prove his conjectures using, in part, the Walnut theorem-prover. We show how his word is related to the infinite Tribonacci word, and we determine both the factor complexity and critical exponent of $\bf B$.
@@ 1 Introduction
In June 2017, Clark Kimberling defined sequence A288462 in the OEIS [16] as follows: it is the infinite fixed point of the inflation rules $00\mapsto 0101$, $1\mapsto 10$, starting with $00$. Because these rules involve a type of substitution more complicated than just a morphism, it is more challenging to analyze. In this note we prove his conjectures using, in part, the Walnut theorem-prover. We also show how his word is related to the infinite Tribonacci word, and we determine both the factor complexity and critical exponent of $\bf B$.
@@
As stated, Kimberling’s original description is perhaps slightly vague, so here is some elaboration. We start with $B_{0}=00$. To find $B_{i+1}$ from $B_{i}$, we do the following: we factor $B_{i}$ into maximal blocks of the form $00$, 0, and $1$; here maximal means we cannot extend a block further to the right or left. Then $B_{i+1}$ is the result of applying the inflation rules $0\mapsto 0$, $1\mapsto 10$, and $00\mapsto 0101$ to $B_{i}$.
@@
Consider $|B_{i}|$, the length of the $i$’th iterate. Let $c_{0}=2$, $c_{1}=4$, $c_{2}=6$, $c_{3}=10$, and define $c_{i}=2c_{i-1}-c_{i-4}$ for $i\geq 4$; this is sequence A288465 in the OEIS. Kimberling conjectured that $|B_{i}|=c_{i}$ for all $i\geq 0$. In this note, we prove Kimberling’s conjecture, as well as his conjectures about the related sequences A288463 and A288464. We also find other properties of the sequence $\bf B$} that link it to the infinite Tribonacci word A080843.
Kimberling indexed the sequence $\bf B$} starting at position $1$. However, for using Walnut, it is easier to index starting at position 0, and this is the convention we use in this paper at the beginning. Later on, in Section 4, we will have to use his indexing.
@@ 2 Kimberling’s conjecture
Let $c_{0}=2$, $c_{1}=4$, $c_{2}=6$, $c_{3}=10$, and define $c_{i}=2c_{i-1}-c_{i-4}$ for $i\geq 4$. Then $|B_{i}|=c_{i}$ for $i\geq 0$.
Proof.
A simple induction now shows that, for $n\geq 1$, the word $B_{i}$ starts with $01$ and contains no occurrences of either $000$ or $111$. Therefore $B_{i}$ can be factorized uniquely as a concatenation of the single initial 0 and blocks of the form $1$, $10$, and $100$. For $x\in\{1,10,100\}$, define $N_{x}(i)$ to be the number of occurrences of the block $x$ in this factorization of $B_{i}$. Then clearly $|B_{i}|=1+N_{1}(i)+2N_{10}(i)+3N_{100}(i)$.
Define a mapping $\xi$ on the blocks $x$ that obeys Kimberling’s inflation rules, and hence sends $1$ to $10$, $10$ to $100$, and $100$ to $100101$. Write $B_{i}=0x_{1}x_{2}\cdots x_{k}$, where each $x_{j}\in\{1,10,100\}$, $1\leq j\leq k$. By checking what happens at the boundaries, we see that $B_{i+1}=0\xi(x_{1})\cdots\xi(x_{k})$.
We now claim that the following relations hold for $i\geq 2$:
$N_{1}(i)$ $=N_{100}(i-1)$ (1)
$N_{10}(i)$ $=N_{100}(i-1)+N_{1}(i-1)$ (2)
$N_{100}(i)$ $=N_{10}(i-1)+N_{100}(i-1).$ (3)
These follow immediately by looking at the image of each block above.
We now show that these identities are enough to directly prove the following (no induction needed!):
$N_{1}(i)$ $=N_{1}(i-1)+N_{1}(i-2)+N_{1}(i-3)$ (4)
$N_{10}(i)$ $=N_{10}(i-1)+N_{10}(i-2)+N_{10}(i-3)$ (5)
$N_{100}(i)$ $=N_{100}(i-1)+N_{100}(i-2)+N_{100}(i-3)$ (6)
@@
Proof of Eq. (4):
[…]
Proof of Eq. (5):
[…]
Proof of Eq. (6):
[…]
We also have
[…]
and substituting into Eq. (7) gives the desired result for $N_{10}(i)$.
Now from above we know that $|B_{i}|-1=N_{1}(i)+2N_{10}(i)+3N_{100}(i)$. Since each of the sequences $(N_{1}(i))_{i}$, $(N_{10}(i))_{i}$, and $(N_{100}(i))_{i}$ on the right-hand side is annihilated by the shift polynomial $X^{3}-X^{2}-X-1$, so is their linear combination $(N_{1}(i)+2N_{10}(i)+3N_{100}(i))_{i}$. And since X1X-1 annihilates the constant sequence $1$, the sequence $(|B_{i}|)_{i}$ is annihilated by the product of X1X-1 and $X^{3}-X^{2}-X-1$, which is $X^{4}-2X^{3}+1$. In other words, $|B_{i}|=2|B_{i-3}|-|B_{i-4}|$ for $i\geq 4$. After comparing the initial conditions $i=0,1,2,3$, it now follows that $|B_{i}|=c_{i}$ for all $i\geq 0$. ∎
@@ 3 Relationship to the Tribonacci word
The infinite sequence $\bf B$ given in A288462 is quite closely related to the celebrated Tribonacci word ${\bf TR}=t_{0}t_{1}t_{2}\cdots=01020100102010102010010201020100102010102010\cdots$, the fixed point of the morphism $0\mapsto 01$, $1\mapsto 02$, $2\mapsto 0$. For more information about $\bf TR$, see, for example, [3].
@@
Kimberling’s sequence $\bf B$} is equal to $0f(\bf TR)$, where $f:0\mapsto 10,\ 1\mapsto 0,\ 2\mapsto 1$.
@@
Let $c_{i}(n)$ denote the number of occurrences of the letter $i$ in the length-$n$ prefix of $\bf TR$. In [14, §10.12], it is shown how to obtain synchronous automata computing the maps $n\mapsto c_{i}(n)$. By “synchronous” we mean that these automata (called c0, c1, c2) take two inputs in parallel, $n$ and $x$, in Tribonacci representation and accept if and only if $x=c_{i}(n)$. See [13] for more about the notion of synchronous automata.
@@
We can now write Walnut code that implements this algorithm. For more about Walnut and its use in combinatorics on words, see [14]. The first 7 lines are taken from [14, §10.12].
@@ 4 Related sequences
Similarly, we can compute $I_{0}(n)$, the index of the $n$’th 0 in $\bf B$, again with $\bf B$ indexed starting at $1$. This is sequence A288463 in the OEIS.
@@
for $n\geq 1$ and some constant $\psi\doteq 1.83$. It turns out that $\psi=1.8392867552\cdots$ is the Tribonacci constant, the unique real zero of the polynomial $X^{3}-X^{2}-X-1$.
@@
    $\lfloor\gamma n\rfloor-1\leq I_{1}(n)\leq\lfloor\gamma n\rfloor+3$.
@@
    Now we show that $1\leq A_{0}(n)-I_{0}(n)\leq 1$:
@@
    For $A_{1}(n)$ we use the code tribb from [14, §10.12]:
@@
    which, by rearranging, implies that
    $3/2\leq I_{1}(n)-\lfloor\gamma n\rfloor\leq 3$
@@
5 Factor complexity and critical exponent
Recall that by factor we mean a contiguous block of letters within a word.
Recall that the factor complexity (aka factor complexity) of a sequence is the function mapping $n$ to the number of distinct blocks of length $n$ appearing in it. A word is called a Rote word [12] if its factor complexity is $2n$ for $n\geq 1$.
@@ 5.1 Bispecial factors of the sequence $\bf B$
In order to determine both the factor complexity and the critical exponent of $\bf B$, the knowledge of bispecial factors in $\bf B$ is essential. The description of bispecial factors and their return words in $\bf TR$ is taken from [6, 8, 7]. The sequence $(b_{n})_{n=0}^{\infty}$ of all non-empty bispecial factors, ordered by length, in the Tribonacci word, satisfies $b_{0}=0$ and for $n\geq 1$,
$b_{n}=\varphi(b_{n-1})0,\quad\text{where}\ \varphi:0\mapsto 01,\ 1\mapsto 02,\ 2\mapsto 0\,.$
@@
By Theorem 2, Kimberling’sequence is equal to $f({\bf TR})$. As $\bf TR$} is fixed by $\varphi$, $\bf B$} is also the image of the Tribonacci sequence under the morphism $\pi=f\circ\varphi$. Instead of $f$ we will work with the morphism $\pi:0\mapsto 100,\ 1\mapsto 101,\ 2\mapsto 10$ because it allows us to derive the form of bispecial factors in $\bf B$ in an easier way.
@@
5.2 Factor complexity of the sequence $\bf B$
@@
By Lemma 9, for each length $n\geq 1$, there are two distinct left special factors in $\bf B$, which confirms that the factor complexity is $2n$ for all $n\geq 1$. ∎
@@ 5.3 Return words to bispecial factors in $\bf B$
    The claim (b) follows immediately from (a).
@@
    $\begin{array}[]{rcl}|v|&\geq&\min\{|\pi(\varphi^{n}(00))|,|\pi(\varphi^{n}(01))|,|\pi(\varphi^{n}(02))|\}=\\ &=&|\pi(\varphi^{n}(02))|=|\pi(\varphi^{n}(0))|+|\pi(\varphi^{n-1}(0))|.\end{array}$